# inkscapepath=svgsubdir

### v1

@@ 8 Conclusion
We have explained the training dynamics of a simplified transformer on the problem of modular addition in terms of two simple phenomena: clustering and alignment. We have provided strong evidence for our model using a particle simulation. Our results are very consistent with the previous findings of Zhong et al. [2023] and Gromov [2023], who showed that circular structures play an important role in the successful generalization of modular addition, and Liu et al. [2022], who studied the generalization properties of similar grid-like structures in the problem of modular addition.
We believe that our findings open up many new research directions in the field of mechanistic interpretability by showing that it is possible to provide mechanistic explanations not only for the inner workings of trained neural networks, but also for the emergence of these inner workings during training. Exciting challenges lie ahead in understanding the training dynamics of more complex models and problems.
Limitations
We have focused on a very simplified setting: solving modular addition with a single-layer transformer with constant attention and 2-dimensional embeddings. Significant challenges lie ahead in understanding the training dynamics of more complex models and problems. Morveover, even in this simplified setting, our model lacks a theoretical framework, being supported only by empirical evidence and a qualitative analysis. Providing mathematical proofs for the forces behind the training dynamics is an important direction for future research.
Broder Impact
We believe that understanding the training dynamics of neural networks can play a crucial role in the development of better optimizers and neural architectures that are more interpretable and safer, but also more accurate and compute-efficient. It is important to apply caution and responsible judgment when utilizing these new techniques.
Compute Resources
All experiments were conducted on a cloud instance with a single NVIDIA T4 GPU, 16 GB of RAM, and 16 vCPUs. For the entire study, the instance was running for 271 hours, but we expect that a replication would take less than 24 hours for both the transformer and the particle simulation.
@@ A.1 Circle detection
To determine whether the embedding vectors form a circle, we simply consider the ratio of the maximum and minimum distances of the embeddings to the origin. If the ratio is below a certain threshold, we consider the embeddings to form a circle. We use a threshold value of 1.2. For our setup, we observe that this simple algorithm is sufficient to detect circles with great accuracy.
@@ A.2 Grid imperfections
We propose a simple measure to quantify the number of grid imperfections inspired by the following fact: in a perfect grid, the vector sums of all pairs of embeddings also form a perfect grid where pair sums are grouped together based on their modular sum.
Our measure works as follows. For each pair of embeddings $(E_{i},E_{j})$, we find the pair of embeddings $(E_{k},E_{l})$ with the closest vector sum. If the modular sum $(i+j)\ \mathrm{mod}\ N$ is different from the modular sum $(k+l)\ \mathrm{mod}\ N$, we consider this pair to be an “imperfection”.
@@ Appendix B Weight decay and the magnitude of weights and biases
Below we visualize the magnitude of the weights and biases of the linear layer at the end of training for various values of weight decay. We run $10$ random initializations for $2000$ epochs with a learning rate of $0.01$. We plot the average absolute value of the weights and biases of the linear layers among the $10$ runs in Figure 6. We observe that weight decay strongly limits the magnitude of the weights and biases, with biases being limited more than weights.
@@
Appendix C Embedding vectors
Below we visualize the embedding vectors that emerge from several random initializations in the trained transformer for various values of weight decay, after training for 2000 epochs with a learning rate of 0.01.
@@ Appendix C Embedding vectors
C.1 $\wWeightdecay\w\mathrm{Weight\ decay}=\w$
@@
Appendix D Particle simulation
Below we visualize several particle arrangements that emerge from different random initializations in the particle simulation for various values of $f_{a}$, after running the simulation for 100 steps. We maintain $N=17,D=2,g_{r}=g_{a}=1$.
@@ Appendix D Particle simulation
Appendix E Classifier visualization
Below we visualize the embedding vectors (left) and the classification function formed by the subsequent layers (right) of several models at the end of training. In the right plot, we also plot the sums of the embedding vectors of all pairs in the training set (“pair sums”). Models were trained for 2000 epochs with a learning rate of 0.01 and a weight decay of 0.6.
@@ Appendix E Classifier visualization
Appendix F Embedding vectors in 3D and 4D
Below we visualize the embedding vectors that emerge from several random initializations in the trained transformer with 3-dimensional and 4-dimensional embeddings. We train for 2000 epochs with a weight decay of 1 and a learning rate of 0.01.
We use three 2D plots to visualize each configuration of embeddings. We project the embeddings onto pairs of dimensions and plot the resulting 2D projections. It is not possible to visualize 3D and 4D embeddings directly, but we can still get a sense of their grid-like or circular structure by examining the 2D projections.
F.1 3-dimensional embeddings

### v2

@@ 7 Conclusion
I have explained the training dynamics of a simplified transformer on the problem of modular addition in terms of two simple phenomena: clustering and alignment. I have provided strong empirical evidence for my model using a particle simulation. I have also provided a qualitative argument for the two phenomena, but further work is needed to fully understand their origin.
@@ A.1 Circle detection
To determine whether the embedding vectors form a circle, I simply consider the ratio of the maximum and minimum distances of the embeddings to the origin. If the ratio is below a certain threshold, I consider the embeddings to form a circle. I use a threshold value of 1.2. For my setup, I observe that this simple algorithm is sufficient to detect circles with great accuracy.
@@ A.2 Grid imperfections
I propose a simple measure to quantify the number of grid imperfections inspired by the following fact: in a perfect grid, the vector sums of all pairs of embeddings also form a perfect grid where pair sums are grouped together based on their modular sum.
My measure works as follows. For each pair of embeddings $(E_{i},E_{j})$, I find the pair of embeddings $(E_{k},E_{l})$ with the closest vector sum. If the modular sum $(i+j)\ \mathrm{mod}\ N$ is different from the modular sum $(k+l)\ \mathrm{mod}\ N$, I consider this pair to be an “imperfection”.
@@ Appendix B Weight decay and the magnitude of weights and biases
Below I visualize the magnitude of the weights and biases of the linear layer at the end of training for various values of weight decay. I run $10$ random initializations for $2000$ epochs with a learning rate of $0.01$. I plot the average absolute value of the weights and biases of the linear layers among the $10$ runs in Figure 6. We can observe that weight decay strongly limits the magnitude of the weights and biases, with biases being limited more than weights.
@@
Appendix C The role of weight decay
$\theta_{t+1}\leftarrow\theta_{t}-\lambda\gamma\theta_{t}$ (4)
Traditionally, weight decay was understood as a form of $L_{2}$ regularization, improving generalization by constraining the network capacity [Goodfellow et al., 2016]. More recently, a new perspective on weight decay has emerged, suggesting that it plays a much more important role during training by changing the training dynamics in a desirable way [Zhang et al., 2018]. In this section, I combine the two perspectives by discussing the impact of weight decay on the training dynamics in my setup.
In Section 4, we observed that weight decay plays an important role in the successful generalization of the trained model. I propose the following explanation: weight decay changes the training dynamics by strengthening the clustering force and by introducing the alignment force. In turn, these forces facilitate the emergence of grids and circles, which are crucial for the generalization performance of the model. Below I discuss exactly how weight decay impacts clustering and alignment.
In the case of clustering, my proposed clustering force (Equations 1 and 2) is highly dependent on a simplified model of the classification function where the decision boundaries are very wide. Weight decay enables this by limiting the magnitude of the weights of the linear layer and output layer, as I show in Appendix B. Without weight decay, classification boundaries could get arbitrarily narrow, leading to a clustering force that is very strong in a narrow region, but practically non-existent elsewhere. In other words, without weight decay, the classifier would overfit to the pair sums and the embeddings would become stuck in a local minimum.
In the case of alignment, weight decay makes the alignment force (Equation 3) possible by limiting the magnitude of the biases of the linear layer, as I show in Appendix B. This forces the ReLU activations of the linear layer to remain very close to the origin, creating the sort of alignment we have seen. Without weight decay, biases could get arbitrarily large and there would not be any specific point around which the alignment could take place.
To the best of my knowledge, this link between regularization and training dynamics has not been previously reported. This opens up interesting new questions for future research, such as providing theoretical proof for this phenomenon or showing how it takes place in other problems and architectures.
Appendix D Embedding vectors
Below I visualize the embedding vectors that emerge from several random initializations in the trained transformer for various values of weight decay, after training for 2000 epochs with a learning rate of 0.01.
@@ Appendix D Embedding vectors
D.1 $\wWeightdecay\w\mathrm{Weight\ decay}=\w$
@@
Appendix E Particle simulation
Below I visualize several particle arrangements that emerge from different random initializations in the particle simulation for various values of $f_{a}$, after running the simulation for 100 steps. I maintain $N=17,D=2,g_{r}=g_{a}=1$.
@@ Appendix E Particle simulation
Appendix F Classifier visualization
Below I visualize the embedding vectors (left) and the classification function formed by the subsequent layers (right) of several models at the end of training. In the right plot, I also plot the sums of the embedding vectors of all pairs in the training set (“pair sums”). Models were trained for 2000 epochs with a learning rate of 0.01 and a weight decay of 0.6.
@@ Appendix F Classifier visualization
Appendix G Embedding vectors in 3D and 4D
Below I visualize the embedding vectors that emerge from several random initializations in the trained transformer with 3-dimensional and 4-dimensional embeddings. I train for 2000 epochs with a weight decay of 1 and a learning rate of 0.01.
I use three 2D plots to visualize each configuration of embeddings. I project the embeddings onto pairs of dimensions and plot the resulting 2D projections. It is not possible to visualize 3D and 4D embeddings directly, but we can still get a sense of their grid-like or circular structure by examining the 2D projections.
G.1 3-dimensional embeddings